suppose that iq scores within a population are normally distributed with a mean 100 and a sd 10 area enclosed under normAL CURVE FOR IQ SCORES ABOVE 130 IS


1
Expert's answer
2022-06-13T15:18:35-0400

We have the normal distribution with "\mu=100" and "\sigma=10."

Let's find the z-score for "X=130." :

"z=\frac{X-\mu}{\sigma}=\frac{130-100}{10}=3."


The area enclosed under normal curve for iq scores above 130 equals the probability that X is greater than 130:

"P(X>130)=P(z>3)=1-(z<3)=1-0.9987=0.0013."


Answer: 0.0013.